OpenStudy (anonymous):

Winnie rented a room at $250 per week. She also had to pay a deposit of $1200. Part A: Write a function to show the total amount of money f(n) that Winnie has to pay for renting the room for n weeks. (4 points) Part B: What is the total rent that Winnie has to pay for using the room for 12 weeks? (4 points) Part C: If the deposit amount decreases by $100, what would be the total rent for 6 weeks? (2 points)

OpenStudy (anonymous):

Okay, lets start. Do you know what to do first for Part A?

OpenStudy (anonymous):

f(n) = 250n + 1200 :D

OpenStudy (anonymous):

BRILLIANT!!! :D

OpenStudy (anonymous):

Part B, do you know what to do here?

OpenStudy (anonymous):

I bet you do. q.q

OpenStudy (anonymous):

Now f(12) = 250(12)+ 1200 f(12) = 3000 + 1200 f(12) = 4200

OpenStudy (anonymous):

Don't mind me, I'm just stalking o^O

OpenStudy (anonymous):

Absolutely perfect. :D

OpenStudy (anonymous):

I've notised @jiji501 xD

OpenStudy (anonymous):

lerl jiji

OpenStudy (anonymous):

C is easy.

OpenStudy (anonymous):

Yeah. :D See how much you can teach someone? :D

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I'm having fun, and it's awesome to see that you are learning it quickly

OpenStudy (anonymous):

So, part C. What is your new function gonna be?

OpenStudy (anonymous):

f(6) = 250(6) + 1300 f(6) = 1500 + 1300 f(6) = 2800

OpenStudy (anonymous):

Decreases. Not increases. woops

OpenStudy (anonymous):

Yeah, derp. <_>

OpenStudy (anonymous):

f(6) = 250(6) + 1100 f(6) = 1500 + 1100 f(6) = 2600

OpenStudy (anonymous):

Indeedly. :D

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Mk lets do the next one

OpenStudy (anonymous):

Okay